i have a word problem and am trying to estimate maximum height and have the equation 0=-4.9t^2+16t+20 and need to solve for t but i cant find a way that works please help
Answer by rothauserc(4718)

-4.9t^2+16t+20 is a parabola that opens downward, so we want the vertex of the parabola
t coordinate for the vertex is -b/2a = -16/(2*-4.9) = 1.632653061
now substitute for t in the equation
max height = -4.9*(1.632653061)^2 +16*1.632653061 + 20 = 33.06122449